# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ 1 Introduction
Random matrix models have long played a central role in two-dimensional (2D) Euclidean quantum gravity, where they provide a direct combinatorial and analytic handle on the sum over random surfaces. In particular, self-adjoint random matrix models [DiFrancesco:1993, DiFrancesco:2004qj] provide a direct handle on the sum over random oriented geometries. Through the dual-graph correspondence, the partition function of this matrix model was shown to be equal to that of Euclidean Dynamical Triangulations (EDT) [Ambjorn:1985az].
@@
EDT is a state-sum formulation of quantum gravity (QG). In this framework, the geometry of spacetime is discretized by a tessellation into intrinsically flat simplices, and the Einstein–Hilbert action is approximated by its Regge version [Regge:1961]. In contrast to quantum field theory (QFT) on a fixed background, in quantum gravity the lattice itself is dynamical, as its structure represents the gravitational degrees of freedom. The partition function is therefore obtained as the continuum limit of a state sum over all tessellations. In the matrix model formulation, the continuum limit is reached by sending the matrix size $N\to\infty$ and tuning the face weights to criticality. Due to the availability of powerful combinatorial and analytic tools [Budd2023] as well as known techniques from high-energy physics, these discrete models become considerably more tractable and their continuum limits can be studied analytically and numerically. For example, Functional Renormalization Group (FRG) techniques provide a systematic way to analyze this limit through the large-$N$ expansion [Eichhorn:2013isa]. Altogether, this motivates the search for analytic frameworks beyond 2D, capable of capturing higher-dimensional generalizations of EDT.
@@
However, in higher dimensions, extensive numerical evidence shows that the continuum limit of EDT does not yield a geometric phase compatible with our universe. Instead, the path integral is dominated by a “crumpled” phase and a tree-like “branched polymer” phase [Ambjorn:1985az]. This shortcoming motivated the development of discrete approaches that enforce a causal structure at the level of the building blocks themselves, most notably Causal Set Theory [Sorkin1987, Surya:2019ndm] and Causal Dynamical Triangulations (CDT) [Ambjorn:1998, Ambjorn:2012jv, Loll:2019rdj]. In $D$-dimensional CDT, simplicial building blocks carry both spacelike and timelike (D1)(D-1)-faces, and gluing rules forbid configurations incompatible with a global time foliation. Crucially, CDT exhibits strong numerical evidence for the appearance of a genuine geometric phase in the continuum limit [Loll:2019rdj]. This motivated the implementation of such causal structures in other QG state-sum formulations, including random matrix models.
@@
In the 2D case, [Benedetti:2008hc] introduced a causal structure using dually weighted matrix models (DWMM) [das_1990, Kazakov:1995gm, Kazakov:1996zm]. In these, the causal structure is imposed at the level of the Feynman rules by assigning different weights to faces of the ribbon graph. Concretely, the action contains interaction terms dressed with fixed “weight matrices” that distinguish timelike from spacelike identifications of edges. In the dual triangulation, these weights control how different types of simplicial building blocks can meet, and thereby implement a combinatorial remnant of causality already at the level of the matrix integral. While this framework successfully reproduces features of CDT in two dimensions, the resulting matrix model is analytically challenging: the presence of nontrivial face weights obscures the usual large-$N$ techniques and complicates the extraction of continuum limits.
@@
By looking for a tool to regain analytic control over dually weighted matrix models, we encounter the intermediate field representation. This technique introduces an auxiliary field that allows one to rewrite quartic (or higher-order) matrix and tensor interactions in simpler forms. In this way, complicated multitrace couplings are effectively replaced by Gaussian integrals over the auxiliary field. Despite its success in several matrix and tensor models [Bonzom2015ColoredTO, Lionni:2016ush, Rivasseau:2010ke, Gurau:2014lua, Nguyen:2014mga, Lionni2016NoteOT, Erbin:2019zug], its application to more general models, such as dually-weighted theories and models with nontrivial propagators and general interactions, has remained largely unexplored.
@@
Building on this foundation, we construct an intermediate field representation for matrix/tensor models with nontrivial propagators and potentials. Furthermore, motivated by the causal matrix models studied in [Benedetti:2008hc, Castro:2020dzt, abranches2025] as well as higher-dimensional analogues in the form of dually-weighted tensor model [Benedetti:2011nn], we consider certain complex matrix and tensor models that include an external rigidity matrix in the propagator. We find that the partition and correlation functions of these complex matrix/tensor models are exactly equivalent to certain self-adjoint two–matrix/tensor models via a dually-weighted intermediate field representations.
@@
The paper is organized as follows. In Section 2, we begin with a summary of the notation and basic identities used throughout the paper. In Section 3, we review dually-weighted matrix models (DWMM), as well as the causal matrix model introduced by Benedetti and Henson [Benedetti:2008hc] and motivate the introduction of an analogous complex matrix model.
@@ 2 Notations and essentials
  • Definition. For $\sigma\in{\rm S}_{n}$ and $M$ a complex $N\times N$ matrix, ${\rm Tr}_{[\sigma]}(M)=\sum_{k_{1},\dots,k_{n}=1}^{N}\prod_{i=1}^{n}M_{k_{i}k_{\sigma(i)}}.$
  • Cycle-type dependence. If $\lambda\vdash n$ corresponds to $\sigma]$, define ${\rm Tr}_{\lambda}(M):=\prod_{j=1}^{\ell(\lambda)}{\rm Tr}(M^{\lambda_{j}})$, ${\rm Tr}_{\lambda}(M)={\rm Tr}_{[\sigma]}(M).$
  • Homogeneity. For $z\in\mathbb{C}$, ${\rm Tr}_{\lambda}(zM)=z^{\|\lambda\|}{\rm Tr}_{\lambda}(M)$ with $\|\lambda\|=\sum_{j}\lambda_{j}=n.$
@@
  • Symmetric group ${\rm S}_{n}$. A partition $\lambda\vdash n$ labels an irreducible representation $\rho^{\lambda}:{\rm S}_{n}\to{\rm GL}(d^{\lambda})$ with character
    $\chi^{\lambda}(\sigma)={\rm Tr}(\rho^{\lambda}(\sigma)),\qquad d^{\lambda}=\chi^{\lambda}(e).$ (2.1)
  • Orthogonality and completeness (discrete). For $\alpha,\beta\in S_{n}$
    $\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi^{\mu}(\sigma)^{*}\chi^{\nu}(\sigma)=\delta^{\mu\nu},\qquad\sum_{\lambda\vdash n}\chi^{\lambda}(\alpha)^{*}\chi^{\lambda}(\beta)=\frac{n!}{|[\alpha]|}\delta_{[\alpha],[\beta]}.$ (2.2)
  • Double-class identity (discrete convolution). $\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi^{\lambda}(\alpha\sigma\beta\sigma^{-1})=\frac{1}{d^{\lambda}}\chi^{\lambda}(\alpha)\chi^{\lambda}(\beta).$
  • Matrix characters of ${\rm GL}(N)$ and $U(N)$. A partition $\lambda\vdash n$ labels an irreducible representation $\rho_{\lambda}:{\rm GL}(N)\to{\rm GL}(d_{\lambda})$ with character $\chi_{\lambda}(M)={\rm Tr}(\rho_{\lambda}(M)).$ If $M$ has eigenvalues $m_{1},\dots,m_{N}$, then $\chi_{\lambda}(M)=s_{\lambda}(m_{i}),$ the Schur polynomial.
@@
    ${\rm Tr}_{[\sigma]}(M)=\sum_{\lambda\vdash n}\chi^{\lambda}(\sigma)\chi_{\lambda}(M),\qquad\chi_{\lambda}(M)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi^{\lambda}(\sigma){\rm Tr}_{[\sigma]}(M).$ (2.5)
@@
  • Identity and transposition operators. For $\bm{a},\bm{b},\bm{c},\bm{d}\in\{1,\dots,N\}^{D}$, ${\mathbb{1}}^{\otimes D}]^{\bm{a}}_{\bm{b}}=\delta^{\bm{a}}_{\bm{b}}$ and $\Sigma^{\bm{a}\bm{c}}_{\bm{b}\bm{d}}=\delta^{\bm{a}}_{\bm{d}}\delta^{\bm{c}}_{\bm{b}}$ .
@@
  • Index contraction. Repeated indices and are summed over. If $\varphi,\psi\in\mathbb{C}_{N}^{\otimes D}$, then
@@
In the literature, “trace invariants for tensors” may be referred to as “tensor invariants” or “bubbles”.
@@
  • Action on index arrays. Let $\bm{k}=(k_{ij})$ be an $n\times D$ array of indices with $k_{ij}\in\{1,\dots,N\}$. The action of $\bm{\alpha}$ on $\bm{k}$ is defined componentwise by $(\bm{\alpha}_{*}\bm{k})_{ij}=k_{\sigma_{j}^{-1}(i)\,j}.$ That is, $\sigma_{j}$ permutes the $i$-index in the $j$-th column.
@@
    where $\bm{k}_{i}$} denotes the $i$-th row of $\bm{k}$}.
@@ 3 Motivation
In this section, we review three types of random matrix models: dually weighted, self–adjoint multi–matrix ones, and complex matrix models. Furthermore, we give a first glance of how complex matrix models with general quadratic terms in the action can realize dually weighted interactions.
@@ 3.1 Dually weighted matrix model
Dually weighted matrix models (DWMM) extend the self-adjoint one–matrix models by allowing independent control of the combinatorial weights of vertices and faces in the ribbon graph expansion. This is implemented using two fixed “background” matrices where their couplings control such weights. In this subsection, we give a brief summary of this.
@@
For a self-adjoint matrix $A\in{\rm H}(N)$, define
@@
The DWMM is defined by
$\mathcal{Z}_{\rm DW}[V]=\int_{{\rm H}(N)}\mathcal{D}A\;e^{-\frac{N}{2}{\rm Tr}(A^{2})-{\rm Tr}\ln(\mathbb{1}^{\otimes 2}-i\,X\otimes(YA))}\;.$ (3.4)
The ribbon graph expansion of this partition function describes discretized surfaces in which both vertex degrees and face lengths can be adjusted independently. This comes from looking at the role of the coupling sequences (3.1) [^4] The original DWMM literature [Kazakov_1996, das_1990] works with the series (3.3) (without the $i$). The factor of $i$ in (3.2) is introduced here only because it arises naturally in the Fourier/intermediate-field representations used later; it does not change the combinatorics. in the potential (3.3): the factor ${\rm Tr}(X^{k})$ provides a weight for kk–valent vertices, and the terms ${\rm Tr}((YA)^{k})$ produce a cycles of length $k$, so that each face of length $k$ is weighted by ${\rm Tr}(Y^{k})$. Thus, in the ribbon graph expansion of this matrix model:
@@
  • a face of length $k$ contributes ${\rm Tr}(Y^{k})=y^{(k)}$.
Therefore, the matrices $X$ and $Y$, independently, determine the vertex and face statistics of the model. This flexibility makes DWMM suitable for encoding local geometric constraints, like the causal matrix model presented in the next subsection 3.2.
@@ 3.2 Causal matrix model
The causal matrix model [Benedetti:2008hc] provides a matrix integral representation of two–dimensional CDT by translating global causal rules into local constraints on ribbon graphs.
@@
In the Feynman expansion of this model, the propagator of $A$  [^5] This is standard physics jargon for $\langle A_{ij}A_{kl}\rangle(g=0)$, in this case. Along this work, we use this term when referring to Feynman expansions. corresponds to spacelike ribbon-graph edges, while that of $B$ (involving $C_{2}$) corresponds to timelike ribbon-graph edges. Ribbon-graph faces containing BB–lines acquire factors ${\rm Tr}(C_{2}^{m})$, hence, by (3.6), contain exactly two BB–lines, reproducing the CDT rigidity condition: each triangle has only two timelike ribbon-graph edges, and their gluing imposes that no spatial topology change is allowed. In the dual triangulation, these rules imply that each dual face contains either two or zero timelike ribbon-graph edges. These local rules enforce the global foliation structure.
@@
Even though (3.7) successfully reproduces a sum over causal triangulations, this success has not translated into analytic control, since the partition function is not exactly computable with known techniques. Motivated by FRG studies [Eichhorn:2013isa, Castro:2020dzt], we propose a change of variables that moves the constant matrix $C_{2}$ from the quartic to the quadratic part of the action in (3.7). Placing the rigidity matrix $C_{2}$ in the propagator is particularly natural, as it modifies the quadratic part of the action and is analogous to introducing a non-trivial dispersion relation in field theory. Somewhat unexpectedly, this change of variables forces us from working with self-adjoint to complex matrix ensembles, since the change of variables is not valid in the self-adjoint model, but is valid in the complex model. Moreover, it leads to the observation that the partition functions of these complex models coincide with those of certain random self-adjoint matrix and tensor models, revealing an equivalence between complex and self-adjoint ensembles.

### v2

@@ 1 Introduction
Random matrix models have long played a central role in two-dimensional (2D) Euclidean quantum gravity, where they provide a direct combinatorial and analytic handle on the sum over random surfaces. In particular, self-adjoint random matrix models [14, 16] provide a direct handle on the sum over random oriented geometries. Through the dual-graph correspondence, the partition function of this matrix model was shown to be equal to that of Euclidean Dynamical Triangulations (EDT) [3].
@@
EDT is a state-sum formulation of quantum gravity (QG). In this framework, the geometry of spacetime is discretized by a tessellation into intrinsically flat simplices, and the Einstein–Hilbert action is approximated by its Regge version [33]. In contrast to quantum field theory (QFT) on a fixed background, in quantum gravity the lattice itself is dynamical, as its structure represents the gravitational degrees of freedom. The partition function is therefore obtained as the continuum limit of a state sum over all tessellations. In the matrix model formulation, the continuum limit is reached by sending the matrix size $N\to\infty$ and tuning the face weights to criticality. Due to the availability of powerful combinatorial and analytic tools [10] as well as known techniques from high-energy physics, these discrete models become considerably more tractable and their continuum limits can be studied analytically and numerically. For example, Functional Renormalization Group (FRG) techniques provide a systematic way to analyze this limit through the large-$N$ expansion [19]. Altogether, this motivates the search for analytic frameworks beyond 2D, capable of capturing higher-dimensional generalizations of EDT.
@@
However, in higher dimensions, extensive numerical evidence shows that the continuum limit of EDT does not yield a geometric phase compatible with our universe. Instead, the path integral is dominated by a “crumpled” phase and a tree-like “branched polymer” phase [3]. This shortcoming motivated the development of discrete approaches that enforce a causal structure at the level of the building blocks themselves, most notably Causal Set Theory [8, 35] and Causal Dynamical Triangulations (CDT) [4, 2, 30]. In $D$-dimensional CDT, simplicial building blocks carry both spacelike and timelike (D1)(D-1)-faces, and gluing rules forbid configurations incompatible with a global time foliation. Crucially, CDT exhibits strong numerical evidence for the appearance of a genuine geometric phase in the continuum limit [30]. This motivated the implementation of such causal structures in other QG state-sum formulations, including random matrix models.
@@
In the 2D case, [7] introduced a causal structure using dually weighted matrix models (DWMMs) [13, 25, 27]. In these, the causal structure is imposed at the level of the Feynman rules by assigning different weights to faces of the ribbon graph. Concretely, the action contains interaction terms dressed with fixed “weight matrices” that distinguish timelike from spacelike identifications of edges. In the dual triangulation, these weights control how different types of simplicial building blocks can meet, and thereby implement a combinatorial remnant of causality already at the level of the matrix integral. While this framework successfully reproduces features of CDT in two dimensions, the resulting models are analytically challenging: the presence of nontrivial face weights obscures the usual large-$N$ techniques and complicates the extraction of continuum limits.
@@
By looking for a tool to regain analytic control over dually weighted matrix models, we encounter the intermediate field representation. This technique introduces an auxiliary field that allows to rewrite quartic (or higher) matrix and tensor interactions in simpler forms. This effectively replaces complicated multitrace couplings by a Gaussian integrals over the auxiliary field. Despite its success in several matrix and tensor models [9, 29, 34, 23, 31, 28, 21, 5], its application to theories with dually-weighted interactions has remained largely unexplored.
@@
Building on this foundation, and motivated by the causal matrix models studied in [7, 12, 1] as well as higher-dimensiomal analogues in the form of dually-weighted tensor models [6], we consider certain complex matrix and tensor models that include an external rigidity matrix in the propagator. We find that the partition and correlation functions of these complex matrix/tensor models are exactly equivalent to certain self-adjoint two–matrix/tensor models via a dually-weighted intermediate field representations. Interestingly, examples of this type of equivalence have been observed in [36, 15]. In this work, we formalize this equivalence for more general models by constructing an intermediate field representation for matrix/tensor models with nontrivial propagators and potentials.
@@
The paper is organized as follows. In Section 2, we begin with a summary of the notation and basic identities used throughout the paper. In Section 3, we review DWMMs, as well as the causal matrix model introduced by Benedetti and Henson [7] and motivate the introduction of an analogous complex matrix model. In Section 4, we summarize our main results.
@@ 2 Notations and essentials
  • Definition. For $\sigma\in{\rm S}_{n}$ and $M$ a complex $N\times N$ matrix, ${\rm Tr}_{[\sigma]}(M)=\sum_{k_{1},\dots,k_{n}=1}^{N}\prod_{i=1}^{n}M_{k_{i}k_{\sigma(i)}}$, where $\sigma]$ denotes the equivalence class of $\sigma$.
  • Cycle-type dependence. Set ${\rm Tr}_{\lambda}(M):=\prod_{j=1}^{n}{\rm Tr}(M^{\lambda_{j}})$, where $\lambda\vdash n$ is the partition corresponding to the equivalence class $\sigma]$. Then ${\rm Tr}_{\lambda}(M)={\rm Tr}_{[\sigma]}(M).$
  • Homogeneity. For $z\in\mathbb{C}$ and $\lambda\vdash n$, ${\rm Tr}_{\lambda}(zM)=z^{n}{\rm Tr}_{\lambda}(M)$.
@@
  • Symmetric group ${\rm S}_{n}$. A partition $\lambda\vdash n$ labels an irreducible representation $\rho_{\lambda}:{\rm S}_{n}\to{\rm GL}(d_{\lambda})$ with character $\chi_{\lambda}(\sigma)={\rm Tr}(\rho_{\lambda}(\sigma))$ and the dimension of the representation is given by $d_{\lambda}=\chi_{\lambda}(e)$.
  • Orthogonality and completeness (discrete). For $\mu,\nu\vdash n$, and $\alpha,\beta\in S_{n}$,
    $\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi_{\mu}(\sigma)^{*}\chi_{\nu}(\sigma)=\delta_{\mu\nu},\qquad\sum_{\lambda\vdash n}\chi_{\lambda}(\alpha)^{*}\chi_{\lambda}(\beta)=\frac{n!}{|[\alpha]|}\delta_{[\alpha][\beta]}.$ (2.1)
  • Double-class identity (discrete convolution). $\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi_{\lambda}(\alpha\sigma\beta\sigma^{-1})=\frac{1}{d_{\lambda}}\chi_{\lambda}(\alpha)\chi_{\lambda}(\beta).$
  • Matrix characters of ${\rm GL}(N)$ and $U(N)$. A partition $\lambda\vdash n$ labels an irreducible representation $\rho_{\lambda}:{\rm GL}(N)\to{\rm GL}(D_{\lambda})$ with character $\chi_{\lambda}(M)={\rm Tr}(\rho_{\lambda}(M))$ and dimension $D_{\lambda}=\chi_{\lambda}(\mathbb{1})$. If $M$ has eigenvalues $m_{1},\dots,m_{N}$, then $\chi_{\lambda}(M)=s_{\lambda}(m_{i}),$ the Schur polynomial.
@@
    ${\rm Tr}_{[\sigma]}(M)=\sum_{\lambda\vdash n}\chi_{\lambda}(\sigma)\chi_{\lambda}(M),\qquad\chi_{\lambda}(M)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi_{\lambda}(\sigma){\rm Tr}_{[\sigma]}(M).$ (2.4)
@@
  • Identity and transposition operators. For $\bm{a},\bm{b},\bm{c},\bm{d}\in\{1,\dots,N\}^{D}$, ${\mathbb{1}}^{\otimes D}]^{\bm{a}}_{\bm{b}}=\delta^{\bm{a}}_{\bm{b}}$ and we define $\Sigma^{\bm{a}\bm{c}}_{\bm{b}\bm{d}}:=\delta^{\bm{a}}_{\bm{d}}\delta^{\bm{c}}_{\bm{b}}$ .
@@
  • Index contraction. Repeated indices are summed over. If $\varphi,\psi\in\mathbb{C}_{N}^{\otimes D}$, then
@@
In the literature and in our paper, “trace invariants for tensors” may be referred to as “tensor invariants” or “bubbles”.
@@
  • Action on index arrays. Let $\bm{\alpha}\in{\rm S}_{n}^{D}$, and let $\bm{k}=(k_{ij})$ be an $n\times D$ array of indices with $k_{ij}\in\{1,\dots,N\}$. The action of $\bm{\alpha}$ on $\bm{k}$ is defined componentwise by $(\bm{\alpha}_{*}\bm{k})_{ij}=k_{\sigma_{j}^{-1}(i)\,j}.$ That is, $\sigma_{j}$ permutes the $i$-index in the $j$-th column.
@@
    where $\bm{k}_{i}$} and $(\bm{\alpha}_{*}\bm{k})_{i}$ respectively denote the $i$-th row of $\bm{k}$} and the $i$-th row of $(\bm{\alpha}_{*}\bm{k})$.
@@ 3 Motivation
In this section, we review three types of random matrix models: dually weighted, self–adjoint multi–matrix, and complex matrix models.
@@ 3.1 Dually weighted matrix model
Dually weighted matrix models (DWMM) extend the self-adjoint one–matrix models by allowing independent control of the combinatorial weights of vertices and faces in the ribbon graph expansion amplitudes. This is implemented using two fixed “background” matrices which control the weights. In this subsection, we give a brief summary of this.
@@
For a self-adjoint matrix $A\in{\rm H}(N)$, specify
@@
The one-matrix DWMM is defined by
$\mathcal{Z}_{\rm DWMM}[V]=\int_{{\rm H}(N)}\mathcal{D}A\;e^{-\frac{N}{2}{\rm Tr}(A^{2})-{\rm Tr}\ln(\mathbb{1}^{\otimes 2}-i\,X\otimes(YA))}\;.$ (3.4)
The ribbon graph expansion of this partition function describes discretized surfaces in which both vertex degrees and face lengths can be adjusted independently. This comes from looking at the role of the coupling sequences (3.1) [^4] The original DWMM literature [26, 13] works with the series (3.3) (without the $i$). The factor of $i$ in (3.2) is introduced here only because it arises naturally in the Fourier/intermediate-field representations used later; it does not change the combinatorics. in the potential (3.3): the factor ${\rm Tr}(X^{k})$ provides a weight for kk–valent vertices, and the terms ${\rm Tr}((YA)^{k})$ produce a cycles of length $k$, so that each face of length $k$ is weighted by ${\rm Tr}(Y^{k})$. Thus, in the ribbon graph expansion of this matrix model:
@@
  • a face of length $k$ contributes ${\rm Tr}(Y^{k})=y^{(k)}$ ,
to the amplitudes. Therefore, the matrices $X$ and $Y$, independently, determine the vertex and face weights of the model. This flexibility makes DWMM suitable for encoding local geometric constraints, like the causal matrix model presented in the next subsection 3.2.
@@ 3.2 Causal matrix model
The causal matrix model [7] provides a matrix integral representation of two–dimensional CDT by translating global causal rules into local constraints on ribbon graphs.
@@
In the Feynman expansion of this model, the propagator of $A$  [^5] This is standard physics jargon for $\langle A_{ij}A_{kl}\rangle(g=0)$, in this case. Along this work, we use this term when refering to Feynman expansions. corresponds to spacelike ribbon-graph edges, while that of $B$ (involving $C_{2}$) corresponds to timelike ribbon-graph edges. In this model, properties of dually-weighted matrix models are incorporated through the matrix $C_{2}$ (see [7] for more details). Ribbon-graph faces containing BB–lines acquire factors ${\rm Tr}(C_{2}^{m})$, hence, by (3.6), contain exactly two BB–lines, reproducing the CDT rigidity condition: each triangle has only two timelike ribbon-graph edges, and their gluing imposes that no spatial topology change is allowed. In the dual triangulation, these rules imply that each dual face contains either two or zero timelike ribbon-graph edges. These local rules enforce the global foliation structure.
@@
Even though (3.7) successfully reproduces a sum over causal triangulations, this success has not translated into analytic control, since the partition function is not exactly computable with existing techniques. Motivated by FRG studies [19, 12], we propose a change of variables that moves the constant matrix $C_{2}$ from the quartic to the quadratic part of the action in (3.7). Placing the rigidity matrix $C_{2}$ in the propagator is particularly natural, as it modifies the quadratic part of the action and is analogous to introducing a non-trivial dispersion relation in field theory. Somewhat unexpectedly, this change of variables forces us from working with self-adjoint to complex matrix ensembles, since the change of variables is not valid in the self-adjoint model, but is valid in the complex model. Moreover, it leads to the observation that the partition functions of these complex models coincide with those of certain random self-adjoint matrix and tensor models, revealing an equivalence between complex and self-adjoint ensembles.